OpenStudy (anonymous):

An airtight box, having a lid of area 80.0 cm^2, is partially evacuated. Atmospheric pressure is 1.01 ´ 10^5 Pa. A force of 108 lb is required to pull the lid off the box. The pressure in the box was:

OpenStudy (irishboy123):

\(F = A .\Delta P\)

OpenStudy (anonymous):

F = 108 lb ? A = 80cm^2 F/A = P ??

OpenStudy (anonymous):

tried convert 108 lb to N

OpenStudy (anonymous):

I got 408.480 N

OpenStudy (anonymous):

then I divide it by 80 cm^2

OpenStudy (anonymous):

I got 480.408/80 = 6.0051

OpenStudy (anonymous):

Idk if this is right pls help

OpenStudy (irishboy123):

http://www.wolframalpha.com/input/?i=solve%3A++%281.01+*+10%5E5%29+-+480.58%2F%2880+*+10%5E%28-4%29%29+ be careful with units :p victorian units to modern metric and cm to m

OpenStudy (anonymous):

May I ask how did you come up with 80/10^4 ?

OpenStudy (anonymous):

that will my last q thankyouuu btw

OpenStudy (irishboy123):

\(1m = 100cm\) \(.01m = 1cm\) \((.01m)^2 = (1cm)^2\)

OpenStudy (anonymous):

so after that I'll use the Pgauge = Ptotal - Patmospheric ? sorry just clearing things up

OpenStudy (irishboy123):

sounds good!! as long as you are happy with it too :-)) ie getting the same answer.